# Aliasing from Galactic Plane Setting in Widefield Radio Interferometry

### v1

@@ Abstract
Measurements with widefield radio interferometers often include the near-infinite gradient between the sky and the horizon. This causes aliasing inherent to the measurement itself, and is purely a consequence of the Fourier basis. For this reason, the horizon is often attenuated by the instrumental beam down to levels deemed inconsequential. However, this effect is enhanced via our own Galactic plane as it sets over the course of a night. We show all-sky simulations of the Galactic plane setting in a radio interferometer in detail for the first time. We then apply these simulations to the Murchison Widefield Array to show that a beam attenuation of 0.1% is not sufficient in some precision science cases. We determine that the noise statistics of the residual data image are drastically more Gaussian with aliasing removal, and explore consequences in simulation for cataloging of extragalactic sources and 21 cm Epoch of Reionization detection via the power spectrum.
@@ 1 Introduction
Foregrounds at the horizon have been seen in Fourier space, creating a ”pitchfork” effect in the Fourier transform along element–element separations known as delay space (Thyagarajan et al., 2015). Calibration accuracy can be improved by filtering the measurements themselves in time as foregrounds move across the night for drift-scan instruments (Charles et al., 2023). Non-interferometric radio observations of the full sky, like in global-signal experiments, see the horizon to a significant degree and thus must account for horizon emission in their models (Bassett et al., 2021). In both delay-space corrections and global-signal models, statistical methods were used to indirectly suppress or observe horizon effects.
We present the most detailed image-space observations and simulations of the setting of the Galactic plane to definitively show the effect of horizon-based aliasing along with its progenitor. Our own Galaxy is an extremely bright source of synchrotron emission, i.e. over 2,000 Jy at 159 MHz at the Galactic Centre (Kriele et al., 2022), and subtends an arc over the whole sky. We show that as it sets over the opaque horizon, even at extremely low levels of beam sensitivity, it creates aliasing that is inherent to the observations themselves and is present throughout the entirety of the image. This has consequences in calibration, image-generation, and noise characterisation for precision science cases in widefield radio interferometry.
@@
: Creation of extragalactic catalogs requires deep images to determine flux density and shape of compact and diffuse sources. Even with advanced deconvolution strategies and a range of LST measurements, the aliasing of the Galactic plane will not fully decorrelate in image-space and will affect the completeness of catalogs in certain regions of sky.
@@
: The redshifted 21-cm line from Hydrogen during the EoR is expected to be extremely faint in comparison to the extragalactic and Galactic foregrounds, however they naturally separate in Fourier space due to their differing spectral signatures. Thus, measurements of this time period are pursued in the power spectrum or similar. Noise-like additions to the power spectrum that do not decorrelate with time will preclude the faint EoR measurement.
In Section 3, we simulate the effects of the setting of the Galactic plane in image space at 180 MHz for interferometric observations, and then narrow our focus to simulated MWA Phase I observations of the one of the coldest patches in the Southern hemisphere. In Section 4, we show how our simulations can mitigate the setting of the Galactic plane in real data and improve the overall noise statistics of the image. We propagate these effects into 1) the cataloging metric of catalog completeness and 2) the power spectrum metric of the 21-cm detection of the EoR in Section 5. Appendix 2 outlines our methodologies and Section 6 summarizes our results.
@@ 2.1 Catalogs and MWA Data
Our example data from the MWA was observed on July 1st, 2014 at frequencies 167–198 MHz. The target field was EoR0, centered on RA 0 h, Dec -27 deg. Data is integrated over 2 sec intervals for a total time of 2 min per observation. Over the course of the night, the EoR0 field is observed in a drift-and-shift mode, where the pointing center is updated every 30 min to keep the field in view, and thus a small range of Local Sidereal Times (LSTs) is covered every night.
@@
[…]
@@
: Near the horizon, we also observe Casseiopia A (Cas A), an extremely bright and extended source with flux density of at $\sim$ 18,500 Jy at 74 MHz (Perley & Butler, 2017). We use the SNR-modelling framework built by Cook et al. (2022) to build a Gaussian component model using high-resolution data from the Very Large Array (VLA).
@@
: We use the catalog from Kriele et al. (2022), natively measured at 159 MHz using the Engineering Development Array 2 (EDA2; Wayth et al. 2021) since it most accurately represents the data. Due to a high gradient in beam sensitivity near the horizon, we transform the spherical harmonics to a high-resolution (1.7’) series of point sources. We only include components within Galactic longitudes of $\pm$ 15 deg to simulate the plane only.
2.2 Simulation and Image Framework
@@
To generate images from both simulated visibilities and observed visibilities, we use Fast Holographic Deconvolution [^8] https://github.com/EoRImaging/FHD (FHD; Sullivan et al. 2012; Barry et al. 2019a). FHD’s resource requirement scales effectively, allowing efficient generation of horizon-to-horizon images. There are three algorithms in FHD which are important in the context of this work:
Calibration
: We must remove instrumental effects through calibration for observed visibilities. To capture per-frequency structure, we use the measured auto-correlations, which are very stable in time for the MWA. Coherent noise is mitigated via an overall scaling to cross-correlations. For a more detailed description of calibration, please see Barry et al. (2019a); Li et al. (2019). Calibration is only required for observed data (Section 4).
Beam Estimation
: The beam sensitivity as a function of sky location is used for both gridding and image contouring. We use a Gaussian decomposition of the MWA beam as described by Barry & Chokshi (2022), which is fully analytic in the image and Fourier plane, mitigating discrete FFT effects.
Image Generation
: We grid each visibility onto a Fourier plane using the Gaussian decomposition estimation as a kernel. We also separately grid the contribution of the beam estimation to generate weights in the Fourier plane. The FFT of these discrete planes then generates the orthogonal projection of the image in uniform weighting. Mask contours are generated at the 1% level of the beam. No image-based cuts or padding is performed in order to keep propagated weights for the power spectrum (Section 5).
@@ 3.1 Unpolarized, unattenuated interferometer
We show the Galactic plane catalog in the top row of Figure 1 for three different LSTs. These LSTs represent the observations for before, during, and after the Galactic plane setting. We purposely choose to simulate only $\pm$ 15 deg away from Galactic plane itself (white contours) to capture the vast majority of its brightness whilst leaving as much as the sky unaltered by simulation as possible.
Figure 1: Simulations of the Galactic plane in uniform weighting. Each column corresponds to a different LST: 344.5 deg (left), 359.8 deg (middle), and 14.6 deg (right). The top row is the input catalog from Kriele et al. (2022) out to $\pm$ 15 deg in Galactic coordinates (white contours), with the Galactic center highlighted (blue star). The middle row is the input catalog as seen by an unpolarized interferometer with the MWA layout and uniform beam ($A(l,m,n)=1$). The bottom row is the input catalog as seen by the MWA in its N–S polarization with beam attenuation out to 1% (white contours). As the Galactic plane sets over the horizon (top row), it aliases over the entirety of the image as seen by an interferometer (middle row). This is attenuated by the instrumental beam, but even a beam value of 1% at the horizon is enough to contaminate most of the image (bottom row).
@@
The level of aliasing depends on the location of the Galactic plane. The aliasing is brightest as the full Galactic plane transitions across the horizon (Figure 1, middle). The Galactic plane aliasing is still significant prior to full set (left), but less in comparison. After the full set (right), the aliasing diminishes significantly. As the Galactic plane sets, the associated aliasing moves, albeit slightly. There can be some decoherence from averaging the LSTs together, and it is in this way that some of the aliasing can be mitigated.
The spatial scales affected by the Galactic plane aliasing translates to a few Mpc to tens of Mpc at a redshift of 6.8. Estimates for EoR bubble sizes at that redshift are in the range of 10 cMpc (Wyithe & Loeb, 2004) up to 100 cMpc (Lin et al., 2016), which could pose extraction issues for EoR tomography if no spectral or time mitigation is in place to disentangle Galactic plane aliasing from the EoR signal.
@@ 3.2 Realistic interferometer
Beam attenuation and polarization will affect the perceived level of Galactic plane aliasing. We show the MWA as an example, which has a large FoV. For an observation at an LST of 359.6 deg (Figure 1, middle), the horizon that the Galactic plane sets over is attenuated to 0.2% by the beam response.
The Galactic plane is so bright, however, that this extreme attenuation does not remove the aliasing, apparent in the bottom row of Figure 1. The aliasing is present at the level of a few mJy/pixel at resolutions of 3.36’ per pixel. As investigated in Section 4, this poses a problem in real data.
@@ 4 Real Data Mitigation
The Galactic plane and its aliasing can be seen in 2 min of data with the MWA given a highly precise data analysis described in Section 2.2.
Figure 2: Removal of the Galactic plane in MWA data. Each column corresponds to a different LST and polarization: 359.8 deg and N–S (left) and 340.9 deg and E–W (right). The top row is a 2-min observation where the extragalactic contribution has been removed, imaged with robust weighting out to the second sidelobe. Galactic plane aliasing is observed as vertical streaking. The middle row is the same observation with the Galactic contribution removed as well, reducing the observed aliasing. The bottom row is the probability distribution function (PDF) and cumulative distribution function (CDF) for the extragalactic-removed image (red histogram, cyan line) and the extragalactic-and-Galactic-removed image (purple histogram, yellow line). The statistics of the image where the Galactic contribution has been removed is more noise-like.
The residual image after removing extragalactic sources is shown in Figure 2 (top row) in Briggs 0 weighting. The Galactic plane is over-saturated in the N–S polarization for a zenith-pointed observation (left), and just barely present in the E–W polarization for an observation two pointings before zenith (right). The aliasing, observed as vertical streaking, is apparent in both polarizations and can easily be seen in the primary lobe for the N–S polarization. Much of the remaining flux density in the image is due to off-plane Galactic diffuse emission, which has not been subtracted in order to avoid contaminating the aliasing contribution.
In contrast, our simulation of the Galactic plane through the MWA is also subtracted from the data in the second row. A drastic reduction in brightness can be seen not only at the location of the Galactic plane, but throughout the entirety of the image due to a successful reconstruction of the aliasing. The aliasing structure is similar in size and shape to the remaining off-planar Galactic diffuse emission, suggesting that mapping the diffuse structure would be exceedingly difficult without aliasing removal.
@@
Subtraction Standard
catalog deviation Skewness Kurtosis
@@
Table 1: Statistics of the residual images depending on the subtraction catalog (E: extragalactic, E & GP: extragalactic and Galactic plane). The first row and second row are the N–S image and E–W image in Figure 2, respectively. Units are in mJy/pixel.
@@ 5 Consequences in Precision Science
We show how the setting of the Galactic plane affects two precision science cases: cataloging of faint, extragalactic radio sources and the 21-cm EoR power spectrum. To generate a base level of understanding, we do not include the effects of calibration.
@@ 5.1 Extragalactic Catalogs
The process of making catalogs of extragalactic sources for population studies results in a well-calibrated, mosiac image with all known corrections applied. This image is then used to perform source finding — algorithms which estimate a background RMS to then pull out peaks which are over a supplied threshold, fitting Gaussians, wavelets, or points to build a model of the structure on the sky. Fainter sources are harder to capture in sources-finding algorithms, and can be subject to Eddington bias, resolution bias, and sensitivity bias.
@@
These bias can be quantified via a completeness metric (i.e. Williams et al. 2016; Hale et al. 2019; Franzen et al. 2019). False sources are injected into the final image which represent a theoretical population of complex sources that follow flux, shape, and positional distributions of known sources. The false sources are then recovered by the source-finding algorithms to quantify the level of flux density that is required to be statistically significant in the image.
@@
We closely follow the methods of Lynch et al. (2021) in calculating the completeness metric, with one exception. We inject the theoretical sources into our simulated images and compare the resulting completeness metrics between image simulations which include the Galactic plane and those which do not. The change in the completeness metric is representative of the effect that the Galactic plane aliasing has on faint-source detection.
We observe a bias in the number of recovered sources of less than $0.5\%$, with changes in integrated flux of only a 0 – 200 mJy for a typical range of LST per pointing with the MWA. Biases are more prevalent in algorithms which use complicated RMS-calculation procedures; for example, cutouts around bright sources (used in Lynch et al. (2021) with pyBDSF) introduces higher bias due to accidental inclusion of aliasing. In general, our calculated bias is low enough to not cause much concern for extragalactic catalog science, and is a good example of how integration helps to reduce the Galactic plane aliasing component.
5.2 21-cm EoR Power Spectrum

### v2

@@ Abstract
Measurements with widefield radio interferometers often include the near-infinite gradient between the sky and the horizon. This causes aliasing inherent to the measurement itself, and is purely a consequence of the Fourier basis. For this reason, the horizon is often attenuated by the instrumental beam down to levels deemed inconsequential. However, this effect is enhanced via our own Galactic plane as it sets over the course of a night. We show all-sky simulations of the Galactic plane setting in a low-frequency radio interferometer in detail for the first time. We then apply these simulations to the Murchison Widefield Array to show that a beam attenuation of 0.1% is not sufficient in some precision science cases. We determine that the noise statistics of the residual data image are drastically more Gaussian with aliasing removal, and explore consequences in simulation for cataloging of extragalactic sources and 21 cm Epoch of Reionization detection via the power spectrum.
@@ 1 Introduction
Foregrounds at the horizon have been seen in Fourier space, creating a “pitchfork” effect in the Fourier transform along element–element separations known as delay space (Thyagarajan et al., 2015). Calibration accuracy can be improved by filtering the measurements themselves in time as foregrounds move across the night for drift-scan instruments (Charles et al., 2023). Noninterferometric radio observations of the full sky, like in global-signal experiments, see the horizon to a significant degree and thus must account for horizon emission in their models (Bassett et al., 2021). In both delay-space corrections and global-signal models, statistical methods were used to indirectly suppress or observe horizon effects.
We present the most detailed image-space observations and simulations of the setting of the Galactic plane to definitively show the effect of horizon-based aliasing along with its progenitor. Our own Galaxy is an extremely bright source of synchrotron emission, e.g., over 2000 Jy at 159 MHz at the Galactic Center (Kriele et al., 2022), and subtends an arc over the whole sky. We show that as it sets over the opaque horizon, even at extremely low levels of beam sensitivity, it creates aliasing that is inherent to the observations themselves and is present throughout the entirety of the image. This has consequences in calibration, image generation, and noise characterization for precision science cases in widefield radio interferometry.
@@
Creation of extragalactic catalogs requires deep images to determine flux density and shape of compact and diffuse sources. Even with advanced deconvolution strategies and a range of local sidereal time measurements, the aliasing of the Galactic plane will not fully decorrelate in image space and could affect the completeness of catalogs in certain regions of sky.
@@
The redshifted 21 cm line from hydrogen during the EoR is expected to be extremely faint in comparison to the extragalactic and Galactic foregrounds; however they naturally separate in Fourier space due to their differing spectral signatures. Thus, measurements of this time period are pursued in the power spectrum or similar. Noise-like additions to the power spectrum that do not decorrelate with time will preclude the faint EoR measurement.
In Section 3, we simulate the effects of the setting of the Galactic plane in image space at 180 MHz for interferometric observations, and then narrow our focus to simulated MWA Phase I observations of one of the coldest patches in the Southern hemisphere. In Section 4, we show how our simulations can mitigate the setting of the Galactic plane in real data and improve the overall noise statistics of the image. We propagate these effects into (1) the cataloging metric of catalog completeness and (2) the power spectrum metric of the 21 cm detection of the EoR in Section 5. Section 2 outlines our methodologies and Section 6 summarizes our results.
@@ 2.1 Catalogs and MWA Data
Our example data from the MWA was observed on 2014 July 1 at frequencies 167–198 MHz. The target field was EoR0, centered on RA 0 h, decl. –27 deg. Data is integrated over 2 s intervals for a total time of 2 min per observation. Over the course of the night, the EoR0 field is observed in a drift-and-shift mode, where the pointing center is updated every 30 min to keep the field in view, and thus a small range of local sidereal times (LSTs) is covered every night.
@@
We use the bespoke LoBES catalog, which was built specifically for the precise requirements of EoR science with the MWA (Lynch et al., 2021). It uses a combination of point-source, Gaussian, and shapelet components to model 80,824 sources in 3069 deg2. LoBES is 70% complete at 10.5 mJy and 90% complete at 32 mJy.
@@
Near the horizon, we also observe Casseiopia A (Cas A), an extremely bright and extended source with flux density of at $\sim$ 18,500 Jy at 74 MHz (Perley & Butler, 2017). We use the supernova remnant modeling framework built by Cook et al. (2022) to build a Gaussian component model using high-resolution data from the Very Large Array (VLA).
@@
We use the catalog from Kriele et al. (2022), natively measured at 159 MHz using the Engineering Development Array 2 (EDA2; Wayth et al. 2021) since it most accurately represents the data. Due to a high gradient in beam sensitivity near the horizon, we transform the spherical harmonics to a high-resolution (1.7) series of point sources. We only include components within Galactic longitudes of $\pm$ 15 deg to simulate the plane only.
2.2 Simulation Framework
@@
These simulated visibilities are then used directly to make images. There are no added effects (e.g., calibration errors or noise) unless otherwise stated. In Section 5.2, we briefly mention the results of expected calibration errors for the 21 cm EoR science case—for this, we use a simple linear least squares solver between the simulated visibilities and reference visibilities to generate per-tile gains. We note that this is only to generate basic understanding, as calibration algorithms are generally much more complex.
To generate images, we use Fast Holographic Deconvolution [^8] https://github.com/EoRImaging/FHD (FHD; Sullivan et al. 2012; Barry et al. 2019a). FHD’s resource requirement scales effectively, allowing efficient generation of horizon-to-horizon images. FHD also includes beam estimation algorithms, which are critical to reducing analysis systematics at the horizon with the MWA.
The beam, used for both gridding visibilities and image contouring, is estimated using a Gaussian decomposition of the MWA sensitivity as described by Barry & Chokshi (2022). This beam is fully analytic in the image and Fourier plane. There are no discrete fast Fourier transform effects in our Fourier plane caused by transforming an instrumentally accurate beam, which is particularly helpful for low sensitivity regions such as the horizon.
We grid each visibility onto a Fourier plane using the Gaussian decomposition estimation as a kernel. We also separately grid the contribution of the beam estimation to generate weights in the Fourier plane. The fast Fourier transform of these discrete planes then generates the orthogonal projection of the image in uniform weighting. Mask contours are generated at the 1% level of the beam. No image-based cuts or padding is performed in order to keep propagated weights for the power spectrum (Section 5.2).
2.3 Data Framework
Our framework for analyzing data in this work, whilst extremely similar to the simulation framework, has some key differences.
We must remove instrumental effects through calibration for observed visibilities. First, we capture per-frequency structure from the measured auto-correlations, which are very stable in time for the MWA, for each tile. Coherent noise is mitigated via an overall scaling to cross-correlations.
We use the aforementioned simulations in Section 2.2 in a linear least squares solver to calculate the per-tile phases, and then fit low-order polynomials to the results. In general, the simulations include over 38,000 sources from the LoBES catalog, a Cas A model, and $\pm$ 15 deg of the Galactic plane.
This combination of auto-correlation and cross-correlation information reduces the dependence on knowledge of the sky structure. For a more detailed description of calibration, please see Barry et al. (2019a); Li et al. (2019).
Gridding and imaging is treated the exact same as simulations in Section 2.2.
@@ 3.1 Unpolarized, unattenuated interferometer
We show the Galactic plane catalog in the top row of Figure 1 for three different LSTs. These LSTs represent the observations for before, during, and after setting of the Galactic center (blue star). We purposely choose to simulate only $\pm$ 15 deg away from Galactic plane itself (white contours) to capture the vast majority of its brightness while leaving as much as the sky unaltered by simulation as possible.
Figure 1: Simulations of the Galactic plane in uniform weighting. Each column corresponds to a different LST: 344.5 deg (left), 359.8 deg (middle), and 14.6 deg (right). The top row is the input catalog from Kriele et al. (2022) out to $\pm$ 15 deg in Galactic coordinates (white contours), with the Galactic center highlighted (blue star). The middle row is the input catalog as seen by an unpolarized interferometer with the MWA layout and uniform beam ($A(l,m,n)=1$). The bottom row is the input catalog as seen by the MWA in N–S polarization out to 1% beam attenuation (white contours). As the brightest part of the Galactic plane (top row, blue star) sets over the horizon, it aliases over the entirety of the image as seen by an interferometer (middle row). This is attenuated by the instrumental beam, but even a beam value of 1% at the horizon is enough to contaminate most of the image (bottom row).
@@
The level of aliasing depends on the location of the Galactic plane and the amount of brightness transitioning across the horizon. The aliasing is brightest as the majority of the Galactic plane and center transitions across the horizon (Figure 1, middle). The Galactic plane aliasing is still significant prior to setting of the Galactic center (left), but less in comparison. After the setting of the Galactic center (right), the aliasing diminishes significantly but is still present due to the remaining brightness of the Galactic plane on the horizon. As the Galactic plane sets, the associated aliasing moves, albeit slightly. There can be some decoherence from averaging the LSTs together, and it is in this way that some of the aliasing can be mitigated.
The spatial scales affected by the Galactic plane aliasing translates to a few megaparsecs to tens of megaparsecs at a redshift of 6.8. Estimates for EoR bubble sizes at that redshift are in the range of 10 cMpc (Wyithe & Loeb, 2004) up to 100 cMpc (Lin et al., 2016), which could pose extraction issues for EoR tomography if no spectral or time mitigation is in place to disentangle Galactic plane aliasing from the EoR signal.
@@ 3.2 Realistic interferometer
Beam attenuation and polarization will affect the perceived level of Galactic plane aliasing. We show the MWA as an example, which has a large field of view (FoV). For an observation at an LST of 359.6 deg (Figure 1, middle), the horizon that the Galactic plane sets over is attenuated to 0.2% by the beam response.
The Galactic plane is so bright, however, that this extreme attenuation does not remove the aliasing, apparent in the bottom row of Figure 1. The aliasing is present at the level of a few mJy pixel-1 at resolutions of 3.36 pixel-1. As investigated in Section 4, this poses a problem in real data.
@@ 4 Real Data Mitigation
Figure 2: Removal of the Galactic plane in MWA data. Each column corresponds to a different LST and polarization: 359.8 deg and N–S (left) and 340.9 deg and E–W (right). The top row is a 2 min observation where the extragalactic contribution has been removed, imaged with robust weighting out to the second sidelobe. Galactic plane aliasing is observed as vertical streaking. The middle row is the same observation with the Galactic contribution removed as well, reducing the observed aliasing. The bottom row is the probability distribution function (PDF) and cumulative distribution function (CDF) for the extragalactic-removed image (red histogram, cyan line) and the extragalactic-and-Galactic-removed image (purple histogram, yellow line). The statistics of the image where the Galactic contribution has been removed is more noise-like.
We now apply our knowledge of the Galactic plane to real data from the MWA. We start with a residual data image after attempting to remove extragalactic sources, shown in Figure 2 (top row) in Briggs 0 weighting. The remaining flux density on the sky is due to (1) mis-subtractions of extragalactic sources, (2) off-plane Galactic diffuse emission, and (3) the Galactic plane and its associated aliasing.
The Galactic plane is over-saturated in the N–S polarization for a zenith-pointed observation (left, top row), and just barely present in the E–W polarization for an observation two pointings before zenith (right, top row). The aliasing, observed as vertical streaking, is apparent in both polarizations and can easily be seen in the primary lobe for the N–S polarization. Much of the remaining flux density in the image is due to off-plane Galactic diffuse emission, which has not been subtracted in order to avoid contaminating the aliasing contribution.
Our simulations of the Galactic plane in Section 3 can also be used to subtract its contribution from the data to a significant degree. This additional subtraction is applied to the second row of Figure 2. A drastic reduction in brightness can be seen not only at the location of the Galactic plane, but also throughout the entirety of the image due to a successful reconstruction of the aliasing. The aliasing structure is similar in size and shape to the remaining off-planar Galactic diffuse emission, suggesting that mapping the diffuse structure would be exceedingly difficult without aliasing removal.
@@
Table 1
Statistics of the Residual Images Depending on the Subtraction Catalog
Subtraction Catalog Standard Deviation Skewness Kurtosis
@@
Note. E: Extragalactic; E & GP: Extragalactic and Galactic plane. The first row and second row are the N–S image and E–W image in Figure 2, respectively. Units are in mJy pixel-1.
@@ 5 Consequences in Precision Science
We show how the setting of the Galactic plane affects two precision science cases: cataloging of faint, extragalactic radio sources and the 21 cm EoR power spectrum.
@@ 5.1 Extragalactic Catalogs
The process of making catalogs of extragalactic sources for population studies results in a well-calibrated, mosiac image with all known corrections applied. This image is then used to perform source finding — algorithms which estimate a background rms to then pull out peaks which are over a supplied threshold, fitting Gaussians, wavelets, or points to build a model of the structure on the sky. Fainter sources are harder to capture in source-finding algorithms, and can be subject to Eddington bias, resolution bias, and sensitivity bias.
@@
These biases can be quantified via a completeness metric (e.g., Williams et al. 2016; Franzen et al. 2019; Hale et al. 2019). False sources are injected into the final image which represent a theoretical population of complex sources that follow flux, shape, and positional distributions of known sources. The false sources are then recovered by the source-finding algorithms to quantify the level of flux density that is required to be statistically significant in the image.
@@
We closely follow the methods of Lynch et al. (2021) in calculating the completeness metric, with one exception. We inject the theoretical sources into our simulated images and compare the resulting completeness metrics between image simulations which include the Galactic plane and those that do not. The change in the completeness metric is representative of the effect that the Galactic plane aliasing has on faint-source detection. We do not include effects of calibration in order to investigate the base level of bias caused by the Galactic plane.
We generate 100 realizations of injected sources and investigate the bias in both number and flux density in their recovery in a simulated, 2 hr image. When the Galactic plane and its aliasing is present in the image, 0.5% more faint sources are recovered. This seemingly counter-intuitive result is due to the rms calculation of the source-finding algorithms subsuming the Galactic plane aliasing.
The Galactic plane aliasing is easily averaged away when large areas of image are used to calculate an underlying rms. However, when techniques like adaptive scaling are used to calculate a background rms, the size scale used can vary. In these cases, Galactic plane aliasing can artificially boost the calculated rms in smaller box sizes, which is subsequently subtracted from the image. This can reveal fainter sources in our idealized simulations. Over 95% of sources which are revealed when the Galactic plane aliasing is subsumed into the rms calculation are less than 200 mJy.
In general, our calculated bias is low enough to not cause much concern for extragalactic catalog science. This is in part due to the help of integration, which can mitigate the Galactic plane aliasing component. Currently, environmental and instrumental variations also cause far more bias (e.g., see Lynch et al. 2021).
5.2 The 21 cm EoR Power Spectrum